g(x)=x^2+3x
h(x)=(x+3)/2
Find (hog)(-1)

Answer by stanbon(75887)

g(x)=x^2+3x
h(x)=(x+3)/2
Find (hog)(-1)
------------------
h[g(-1)]
=h[(-1)^2+3(-1)]
=h[-2]
=(-2+3)/2
=1/2
Cheers,
Stan H.